# Nonlinear Kinematics of Recursive Origami Inspired by Spidron

### v1

@@ Abstract
Non-periodic folding of periodic crease patterns paves the way to novel nonlinear phenomena that cannot be feasible through periodic folding. This paper focuses on the non-periodic folding of recursive crease patterns generalized from Spidron. Although it is known that Spidron has a $1$-DOF isotropic rigid folding motion, its general kinematics and dependence on the crease pattern remain unclear. Using the kinematics of a single unit cell of Spidron and the recursive construction of the folded state of multiple unit cells, we consider the folding of Spidron that is not necessarily isotropic. We found that as the number of unit cells increases, the non-periodic folding is restricted and the isotropic folding becomes dominant. Then, we analyze the three kinds of isotropic folding modes by constructing $1$-dimensional dynamical systems governing each of them. We show that these systems can possess different recursive natures depending on folding modes even in an identical crease pattern. Furthermore, we show their novel nonlinear nature, including the period-doubling cascade leading to the emergence of chaos.
@@ 1 Introduction
2 Crease Pattern of Generalized Spidron
@@
[…]
@@
where we fix the scale of 0-th ring by $l_{0}=1$. In the case of $\theta=\pi/4$ in which $r^{1}_{t}=r^{2}_{t}$ holds, the crease pattern has the additional mirror symmetries. We call such a crease pattern symmetric.
@@ 3 Kinematics of a single ring
This section considers the kinematics of a single ring of the generalized Spidron. The kinematics of rings are equivalent independent of their index $t$ because their crease patterns are scaled versions of each other. Hence, we focus on the kinematics of the outermost 0-th ring. For the convenience of the notation, we omit the subscript 0 for quantities such as edge lengths or vertices. In addition, note that we allow self-intersections of facets to occur. After parameterizing the folding of a single ring, we will consider its configuration space. Then, we will consider folded states having a certain symmetry, which we call isotropic.
@@ 3.1 Parameterization
Figure 2: Construction of a folded state of a ring parameterized by $(\rho^{1},\rho^{2},\phi^{1},\phi^{2})$: (1) Project the outer boundary $\mathbf{p}^{0}\mathbf{p}^{1}\mathbf{p}^{2}\mathbf{p}^{3}$ specified by $\rho^{1}$ and $\rho^{2}$ to xyxy plane. (2) Scale the projected boundary as it becomes the unit rhombus. (3) Rotate triangles lying on xy plane which has the metric prescribed by $\phi^{1}$ and $\phi^{2}$. (4) Rotate triangles around $\mathbf{v}^{i}$ so that their base edges become parallel with the edges of the outer boundary in (1). (5) Translate triangles so that their vertices match with $\mathbf{p}^{i}$. (6) If $|\mathbf{q}^{0}\mathbf{q}^{1}|=|\mathbf{q}^{2}\mathbf{q}^{3}|=s$ and $|\mathbf{q}^{1}\mathbf{q}^{2}|=|\mathbf{q}^{3}\mathbf{q}^{0}|=s$, $(\rho^{1},\rho^{2},\phi^{1},\phi^{2})$ represents a valid folded state of 0-th ring.
@@ 3.2 General Folding
The configuration space of a single ring is the set of all its possible folded states and can be parameterized as the solution space of Equation (5) (Figure 3 (A)). As a simple calculation that the number of variables minus the number of constraints equals $2$ shows, the numerically calculated solution space of Equation (5) forms a $2$-dimensional surface in the parameter space; therefore, a single ring of Spidron behaves as a $2$-DOF mechanism. Note that the folded state of the ring is not uniquely determined only from the configuration of its outer boundary because Equation (5) can have multiple solutions of $\phi^{1},\phi^{2}$ for some values of $\rho^{1},\rho^{2}$ (Figure 3 (B) and (C)).
In general, configuration spaces of origami have symmetry, i.e., a transformation that takes a solution to another solution, representing the reversal of the front and back of a paper. In our case, this symmetry can be represented as $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})$ $\mapsto(-\overline{\rho}^{1},-\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})$, where $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})$ is a solution of Equation (5). Here, the $4$-fold symmetry of the crease pattern and the definition of the parameters give the solution space of Equation (5) an additional symmetry, which is represented as $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})\mapsto(\overline{\rho}^{2},\overline{\rho}^{1},$$\overline{\phi}^{2},$$\overline{\phi}^{1})$. Furthermore, if the crease pattern is symmetric as defined in Section 2, the solution space gets additional mirror symmetries: $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})\mapsto(\overline{\rho}^{2},\overline{\rho}^{1},$$\overline{\phi}^{1},$$\overline{\phi}^{2})$ and $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})\mapsto(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{2},$$\overline{\phi}^{1})$.
@@ 3.3 Isotropic Folding
Geometric constraints in isotropic folding can be obtained by substituting $\rho^{1}=-\rho^{2}=\rho$ and $\phi^{1}=-\phi^{2}=\phi$ into Equation (5). Then, because of the $4$-fold symmetry of the crease pattern, the two equations fall into a single equation, which is quartic in terms of $\tau:=\tan{\phi}$:
@@
We observed the existence of two isotropic folding modes also in the non-symmetric crease patterns (Figure 5 (A)). These modes may be characterized using the notion of pop-up/down assignments on interior vertices [Chen and Santangelo 18]. Intuitively, we can distinguish the two modes by the rotation direction of the inner boundary relative to the outer boundary viewed from the top (Figure 5 (B) and (C)). If the inner boundary is rotated anti-clockwise/clockwise, we call these modes pro-/anti-rotation mode, based on the fact that the inner boundary is rotated anti-clockwise relative to the outer boundary on non-symmetric crease patterns. In the following, we write $f_{\mathrm{pro}}$ and $f_{\mathrm{anti}}$ for the mappings which map $\rho^{\mathrm{out}}$ to $\rho^{\mathrm{in}}$ representing the pro- and anti-rotation mode, respectively. In addition, let $\underline{x},\overline{x}]$ denote the domain of these mappings.
Figure 5: (A) Plots of $(\rho^{\mathrm{out}},\rho^{\mathrm{in}})$ representing isotropic folding motions of a ring for the crease patterns in Figure 1 (A). The rightmost plots corresponding to the symmetric crease patterns are mirror symmetric about the horizontal axis. In each diagram, the boundary of $\underline{x},\overline{x}]\times[\underline{x},\overline{x}]$ is indicated, where $\underline{x},\overline{x}]$ is the domain of $\rho^{\mathrm{out}}$ in which $\rho^{\mathrm{in}}$ exists. (B)–(C) Pro-rotation, and anti-rotation mode in $(s,\theta)=(1/2,\pi/12)$. The points representing each of the folded states are plotted in (A).
@@ 4 Kinematics of Multiple Rings
This section considers the kinematics of the Spidron with multiple rings based on the following recurrence construction. Given a folded state of the $t$-th ring, its inner boundary serves as the outer boundary of the $(t+1)$-st ring. Then, we can obtain a folded state of the $(t+1)$-st ring that is compatible with its outer boundary by solving the geometric constraints. The entire folded state can be constructed by carrying out the above procedure iteratively.
@@ 4.1 General Folding: Convergence to Isotropic Folding
Once we set values of $\rho^{1}_{0}$ and $\rho^{2}_{0}$, we can get $\{(\rho^{1}_{t},$ $\rho^{2}_{t},$ $\phi^{1}_{t},$ $\phi^{2}_{t})\}_{t=0,\dots,T-1}$ representing a folded state of the Spidron with $T$ rings through solving Equation (5) iteratively. Then, we can project the configuration space of Spidron with $T$ rings to that of the 0-th ring by projecting a sequence $\{(\rho^{1}_{t},\rho^{2}_{t},\phi^{1}_{t},\phi^{2}_{t})\}_{t=0,\dots,T-1}$ to $(\rho^{1}_{0},\rho^{2}_{0},\phi^{1}_{0})$ space (Figure 6). Note that some sequences can be projected onto the identical point because there can be multiple configurations compatible with the given outer boundary as we saw in Section 3.
Figure 6: Configuration space of multiple rings and recursive construction of a folded state under the fixed crease pattern $(s,\theta)=(0.45,\pi/4)$. (A)–(C) Configuration space of single, double, and triple rings, respectively. The black curve on the gray surface shows the isotropic folding modes of a single ring. (D) Recursive construction from a folded state represented by $(\rho^{1}_{0},\rho^{2}_{0},\phi^{1}_{0},\phi^{2}_{0})=(1.847,-1.860,1.140,-0.768)$ plotted as black circles in (A)–(C). For this 0-th folded state, there are two different folded states for the $1$-st ring. Then, there is no folded state of the 22-nd ring compatible with either of the $1$-st ones.
@@ 4.2 Isotropic Folding
Recurrence behaviors under the prescribed mode assignment can be qualitatively/quantitatively understood using a dynamical system model. Owing to the mode-assignment for each ring, we can uniquely determine $\rho^{\mathrm{in}}_{t}=\rho^{\mathrm{out}}_{t+1}$ from $\rho^{\mathrm{out}}_{t}$. In particular for the pro- and anti-rotation modes, the mapping $\rho^{\mathrm{out}}_{t}\mapsto\rho^{\mathrm{out}}_{t+1}$ is given by $f_{\mathrm{pro}}$ and $f_{\mathrm{anti}}$ for every $t$. Thus, $f_{\mathrm{pro}}$ and $f_{\mathrm{anti}}$ serve as the recurrence relation in the pro- and anti-rotation mode, respectively, which are $1$-dimensional discrete dynamical systems. For the pleats mode, we can define its dynamical system $f_{\mathrm{pleats}}$ as $f_{\mathrm{anti}}\circ f_{\mathrm{pro}}:\rho^{\mathrm{out}}_{t}\mapsto\rho^{\mathrm{out}}_{t+2}$. Note that if $f_{\sigma}([\underline{x},\overline{x}])\subseteq[\underline{x},\overline{x}]$ and $\rho_{0}^{\mathrm{out}}\in[\underline{x},\overline{x}]$, $\rho^{\mathrm{out}}_{t}$ exists for all $t$, which ensures the existence of a corresponding folded state ($\sigma\in\{\mathrm{pro},\mathrm{anti},\mathrm{pleats}\}$). This condition can be judged visually by checking whether the graph of $f_{\sigma}$ is included in a square representing $\overline{x},\underline{x}]\times[\overline{x},\underline{x}]$ or not (Figure 5).
@@
Even if the crease patterns are identical, recurrence behaviors can change dramatically depending on mode assignments because different modes are governed by different dynamical systems. If we increase the number of rings infinitely, the behavior of an orbit $\{\rho^{\mathrm{out}}_{t}\}_{t=0,1,\dots}$ in the limit $t\rightarrow\infty$ is significant. One typical behavior is the convergence to a stable fixed point satisfying $f_{\sigma}(\rho^{\mathrm{out}})=\rho^{\mathrm{out}}$ and $|f^{\prime}_{\sigma}(\rho^{\mathrm{out}})|<1$. Note that the folded state corresponding to the orbit initiated from a fixed point has a self-similarity same as the crease pattern. Graphically, we can detect fixed points by taking the intersection of the graphs of $\rho^{\mathrm{in}}=\rho^{\mathrm{out}}$ and $f_{\sigma}(\rho^{\mathrm{out}})$ (Figure 5). For example, $\rho^{\mathrm{out}}=0$ corresponding to the developed state is always a fixed point of $f_{\sigma}$ independent of crease pattern and $\sigma$, and the orbits in Figure 7 (B) and (C) converge to this point. Such convergence to a developed state has been reported in previous work [Szilassi 04]. However, in Figure 7 (A), the orbits converge to the non-trivial fixed point corresponding to the finitely folded state rather than the flat state that is unstable there. Such non-trivial behavior has not been reported in previous works, and it becomes possible because we generalized the crease pattern.

### v2

# Nonlinear Kinematics of Recursive Origami Inspired by the Spidron

@@ Abstract
Non-periodic folding of periodic crease patterns paves the way to novel nonlinear phenomena that cannot be feasible through periodic folding. This paper focuses on the non-periodic folding of recursive crease patterns generalized from Spidron. Although it is known that Spidron has a $1$-DOF isotropic rigid folding motion, its general kinematics and dependence on the crease pattern remain unclear. Using the kinematics of a single unit cell of the Spidron and the recursive construction of the folded state of multiple unit cells, we consider the folding of the Spidron that is not necessarily isotropic. We found that as the number of unit cells increases, the non-periodic folding is restricted and the isotropic folding becomes dominant. Then, we analyze the three kinds of isotropic folding modes by constructing $1$-dimensional dynamical systems governing each of them. We show that these systems can possess different recursive natures depending on folding modes even in an identical crease pattern. Furthermore, we show their novel nonlinear nature, including the period-doubling cascade leading to the emergence of chaos.
@@ 1 Introduction
2 Crease Pattern of the Generalized Spidron
@@
In the following, we index rings by $t=0,1,\dots$ towards the inside, and write $p^{i}_{t}$, and $q^{i}_{t}=p^{i}_{t+1}$ for the vertices in their outer, and inner squares, respectively ($i\equiv 0,1,2,3\pmod{4}$). Then, the lengths of creases $l_{t}:=|p^{i}_{t}p^{i+1}_{t}|$, $r^{1}_{t}:=|q^{i}_{t}p^{i}_{t}|$, and $r^{2}_{t}:=|q^{i}_{t}p^{i+1}_{t}|$ can be written down using $s$ and $\theta$ as follows:
@@
where we fix the scale of 00-th ring by $l_{0}=1$. In the case of $\theta=\pi/4$ in which $r^{1}_{t}=r^{2}_{t}$ holds, the crease pattern has the additional mirror symmetries. We call such a crease pattern symmetric.
@@ 3 Kinematics of a single ring
This section considers the kinematics of a single ring of the generalized Spidron. The kinematics of rings are equivalent independent of their index $t$ because their crease patterns are scaled versions of each other. Hence, we focus on the kinematics of the outermost 00-th ring. For the convenience of the notation, we omit the subscript 00 for quantities such as edge lengths or vertices. In addition, note that we allow self-intersections of facets to occur. After parameterizing the folding of a single ring, we will consider its configuration space. Then, we will consider folded states having a certain symmetry, which we call isotropic. Note that we used Mathematica to derive and numerically solve the complex equations in this section.
@@ 3.1 Parameterization
Figure 2: Construction of a folded state of a ring parameterized by $(\rho^{1},\rho^{2},\phi^{1},\phi^{2})$: (1) Project the outer boundary $\mathbf{p}^{0}\mathbf{p}^{1}\mathbf{p}^{2}\mathbf{p}^{3}$ specified by $\rho^{1}$ and $\rho^{2}$ to $xy$ plane. (2) Scale the projected boundary as it becomes the unit rhombus. (3) Rotate triangles lying on xy plane which has the metric prescribed by $\phi^{1}$ and $\phi^{2}$. (4) Rotate triangles around $\mathbf{v}^{i}$ so that their base edges become parallel with the edges of the outer boundary in (1). (5) Translate triangles so that their vertices match with $\mathbf{p}^{i}$. (6) If $|\mathbf{q}^{0}\mathbf{q}^{1}|=|\mathbf{q}^{2}\mathbf{q}^{3}|=s$ and $|\mathbf{q}^{1}\mathbf{q}^{2}|=|\mathbf{q}^{3}\mathbf{q}^{0}|=s$, $(\rho^{1},\rho^{2},\phi^{1},\phi^{2})$ represents a valid folded state of 00-th ring.
@@ 3.2 General Folding
The configuration space of a single ring is the set of all its possible folded states and can be parameterized as the solution space of Equation (5) (Figure 3 (A)). As a simple calculation that the number of variables minus the number of constraints equals $2$ shows, the numerically calculated solution space of Equation (5) forms a $2$-dimensional surface in the parameter space; therefore, a single ring of the Spidron behaves as a $2$-DOF mechanism. Note that the folded state of the ring is not uniquely determined only from the configuration of its outer boundary because Equation (5) can have multiple solutions of $\phi^{1},\phi^{2}$ for some values of $\rho^{1},\rho^{2}$ (Figure 3 (B) and (C)).
In general, configuration spaces of origami have symmetry, i.e., a transformation that takes a solution to another solution, representing the reversal of the front and back of a paper. In our case, this symmetry can be represented as $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})$ $\mapsto(-\overline{\rho}^{1},-\overline{\rho}^{2},$$-\overline{\phi}^{1},$$-\overline{\phi}^{2})$, where $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})$ is a solution of Equation (5). Here, the $4$-fold symmetry of the crease pattern and the definition of the parameters give the solution space of Equation (5) an additional symmetry, which is represented as $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})\mapsto(\overline{\rho}^{2},\overline{\rho}^{1},$$\overline{\phi}^{2},$$\overline{\phi}^{1})$. Furthermore, if the crease pattern is symmetric as defined in Section 2, the solution space gets additional mirror symmetries: $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})\mapsto(\overline{\rho}^{2},\overline{\rho}^{1},$$\overline{\phi}^{1},$$\overline{\phi}^{2})$ and $(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{1},$$\overline{\phi}^{2})\mapsto(\overline{\rho}^{1},\overline{\rho}^{2},$$\overline{\phi}^{2},$$\overline{\phi}^{1})$.
@@ 3.3 Isotropic Folding
Geometric constraints in isotropic folding can be obtained by substituting $\rho^{1}=-\rho^{2}=\rho$ and $\phi^{1}=-\phi^{2}=\phi$ into Equation (5). Then, because of the $4$-fold symmetry of the crease pattern, the two equations fall into a single equation, which is quartic in terms of $\tau:=\tan{(\phi/2)}$:
@@
We observed the existence of two isotropic folding modes also in the non-symmetric crease patterns (Figure 5 (A)). These modes may be characterized using the notion of pop-up/down assignments on interior vertices [Chen and Santangelo 18]. Intuitively, we can distinguish the two modes by the rotation direction of the inner boundary relative to the outer boundary viewed from the top (Figure 5 (B) and (C)). If the inner boundary is rotated anti-clockwise/clockwise, we call these modes pro-/anti-rotation mode, based on the fact that the inner boundary is rotated anti-clockwise relative to the outer boundary on non-symmetric crease patterns. In the following, we write $f_{\mathrm{pro}}$ and $f_{\mathrm{anti}}$ for the mappings which map $\rho^{\mathrm{out}}$ to $\rho^{\mathrm{in}}$ representing the pro- and anti-rotation mode, respectively. In addition, let $[\underline{x},\overline{x}]$ denote the domain of these mappings.
Figure 5: (A) Plots of $(\rho^{\mathrm{out}},\rho^{\mathrm{in}})$ representing isotropic folding motions of a ring for the crease patterns in Figure 1 (A). The rightmost plots corresponding to the symmetric crease patterns are mirror symmetric about the horizontal axis. In each diagram, the boundary of $[\underline{x},\overline{x}]\times[\underline{x},\overline{x}]$ is indicated, where $[\underline{x},\overline{x}]$ is the domain of $\rho^{\mathrm{out}}$ in which $\rho^{\mathrm{in}}$ exists. (B)–(C) Pro-rotation, and anti-rotation mode in $(s,\theta)=(1/2,\pi/12)$. The points representing each of the folded states are plotted in (A).
@@ 4 Kinematics of Multiple Rings
This section considers the kinematics of the Spidron with multiple rings based on the following recurrence construction. Given a folded state of the $t$-th ring, its inner boundary serves as the outer boundary of the $(t+1)$-st ring. Then, we can obtain a folded state of the $(t+1)$-st ring that is compatible with its outer boundary by solving the geometric constraints represented by Equation (5) or (6). The entire folded state can be constructed by carrying out the above procedure iteratively. Same as Section 3, we used Mathematica to numerically solve the equations.
@@ 4.1 General Folding: Convergence to Isotropic Folding
Once we set values of $\rho^{1}_{0}$ and $\rho^{2}_{0}$, we can get $\{(\rho^{1}_{t},$ $\rho^{2}_{t},$ $\phi^{1}_{t},$ $\phi^{2}_{t})\}_{t=0,\dots,T-1}$ representing a folded state of the Spidron with $T$ rings through solving Equation (5) iteratively. Then, we can project the configuration space of the Spidron with $T$ rings to that of the 00-th ring by projecting a sequence $\{(\rho^{1}_{t},\rho^{2}_{t},\phi^{1}_{t},\phi^{2}_{t})\}_{t=0,\dots,T-1}$ to $(\rho^{1}_{0},\rho^{2}_{0},\phi^{1}_{0})$ space (Figure 6). Note that some sequences can be projected onto the identical point because there can be multiple configurations compatible with the given outer boundary as we saw in Section 3.
Figure 6: Configuration space of multiple rings and recursive construction of a folded state under the fixed crease pattern $(s,\theta)=(0.45,\pi/4)$. (A)–(C) Configuration space of single, double, and triple rings, respectively. The black curve on the gray surface shows the isotropic folding modes of a single ring. (D) Recursive construction from a folded state represented by $(\rho^{1}_{0},\rho^{2}_{0},\phi^{1}_{0},\phi^{2}_{0})=(1.847,-1.860,1.140,-0.768)$ plotted as black circles in (A)–(C). For this 00-th folded state, there are two different folded states for the $1$-st ring. Then, there is no folded state of the $2$-nd ring compatible with either of the $1$-st ones.
@@ 4.2 Isotropic Folding
Recurrence behaviors under the prescribed mode assignment can be qualitatively/quantitatively understood using a dynamical system model. Owing to the mode-assignment for each ring, we can uniquely determine $\rho^{\mathrm{in}}_{t}=\rho^{\mathrm{out}}_{t+1}$ from $\rho^{\mathrm{out}}_{t}$. In particular for the pro- and anti-rotation modes, the mapping $\rho^{\mathrm{out}}_{t}\mapsto\rho^{\mathrm{out}}_{t+1}$ is given by $f_{\mathrm{pro}}$ and $f_{\mathrm{anti}}$ for every $t$. Thus, $f_{\mathrm{pro}}$ and $f_{\mathrm{anti}}$ serve as the recurrence relation in the pro- and anti-rotation mode, respectively, which are $1$-dimensional discrete dynamical systems. For the pleats mode, we can define its dynamical system $f_{\mathrm{pleats}}$ as $f_{\mathrm{anti}}\circ f_{\mathrm{pro}}:\rho^{\mathrm{out}}_{t}\mapsto\rho^{\mathrm{out}}_{t+2}$. Note that if $f_{\sigma}([\underline{x},\overline{x}])\subseteq[\underline{x},\overline{x}]$ and $\rho_{0}^{\mathrm{out}}\in[\underline{x},\overline{x}]$, $\rho^{\mathrm{out}}_{t}$ exists for all $t$, which ensures the existence of a corresponding folded state ($\sigma\in\{\mathrm{pro},\mathrm{anti},\mathrm{pleats}\}$). This condition can be judged visually by checking whether the graph of $f_{\sigma}$ is included in a square representing $[\overline{x},\underline{x}]\times[\overline{x},\underline{x}]$ or not (Figure 5).
@@
Even if the crease patterns are identical, recurrence behaviors can change dramatically depending on mode assignments because different modes are governed by different dynamical systems. If we increase the number of rings infinitely, the behavior of an orbit $\{\rho^{\mathrm{out}}_{t}\}_{t=0,1,\dots}$ in the limit $t\rightarrow\infty$ is significant. One typical behavior is the convergence to a stable fixed point satisfying $f_{\sigma}(\rho^{\mathrm{out}})=\rho^{\mathrm{out}}$ and $|f^{\prime}_{\sigma}(\rho^{\mathrm{out}})|<1$. Note that the folded state corresponding to an orbit initiated from a fixed point exhibits a self-similarity whose scaling ratio is the same as the crease pattern. This self-similarity is because the orbit of a fixed point forms a constant sequence, meaning that all the rings in the folded state are similar. Graphically, we can detect fixed points by taking the intersection of the graphs of $\rho^{\mathrm{in}}=\rho^{\mathrm{out}}$ and $f_{\sigma}(\rho^{\mathrm{out}})$ (Figure 5). For example, $\rho^{\mathrm{out}}=0$ corresponding to the developed state is always a fixed point of $f_{\sigma}$ independent of crease pattern and $\sigma$, and the orbits in Figure 7 (B) and (C) converge to this point. Such convergence to a developed state has been reported in previous work [Szilassi 04]. However, in Figure 7 (A), the orbits converge to the non-trivial fixed point corresponding to the finitely folded state rather than the flat state that is unstable there. Such non-trivial behavior has not been reported in previous works, and it becomes possible because we generalized the crease pattern.